jurabib FAQ

Jens Berger

                 jb@jurabib.org

February 3, 2004

This is the jurabib-FAQ. It will grow from time to time. If you have suggestions, please send a question and (if possible) the answer to jb@jurabib.org.

Q: How can I format the number and/or the series field ?

A: jurabib provides a markup command for most entry types: \numberandseries. This command contains two arguments: #1 represents the number, #2 the series field.

Let’s take a look at the default definition. It will format

, <series> <number>
\DeclareRobustCommand{\numberandseries}[2]{%
      \unskip\unskip,\space\bibsnfont{#2}%
      \ifthenelse{\equal{#1}{}}{}{~#1}%
}%

If you want to have e.g.

(=<series>. <number>.)

try the following:

\DeclareRobustCommand{\numberandseries}[2]{%
      \unskip\unskip%, <= comma removed
      \space\bibsnfont{(=#2}%
%                      ^opening ...
      \ifthenelse{\equal{#1}{}}{)}{~#1)}%
%                               ^^... closing, if number is given
%                               | ... closing, if number is missing
}%